# Euclid Quick Data Release (Q1) point-like X-ray selected sources

### v1

@@ 7.3 AGN type
In Fig. 18, we plot the respective -band magnitudes as a function of the soft X-ray fluxes for all CTPs classified as extragalactic. We add the —X/O— = 1 lines (Maccacaro et al. 1988), which highlight the typical AGN locus (see also e.g., Salvato et al. 2011; Civano et al. 2012), defined as
where is the X-ray flux in a given energy range, and correspond to the magnitude and a constant which both depend on the specific filter used in the optical observations. For we find . To further assess the type of our sources, we colour-code the plots in the bottom row of Fig. 18 by the X-ray luminosity and the top row by hardness ratio (Park et al. 2006; Cappelluti et al. 2007; Salvato et al. 2008), defined as
where a negative HR value indicates (type I) unobscured sources, whereas a positive HR value suggests harder, likely obscured (type II) sources. Therefore, we include fluxes from the hard X-ray band where available. Specifically, we consider fluxes from the 2.0 – 4.5 keV range for 4XMM DR14, the 2.0 – 7.0 keV range for Chandra, and the 2.3 – 5.0 keV energy range for eROSITA DR1. This ensures as much consistency as possible across the different surveys, allowing us to utilise similar energy bands for comparison and analysis.
Since the markers in Fig. 18 are plotted in order of increasing hardness ratio, many sources in the background are being covered up and predominantly fall within the X/O boundaries, classifying them as bright, unobscured type I sources, with different populations visible for every X-ray survey. This trend is most pronounced for eROSITA, where the distribution is more concentrated within the boundaries. A similar pattern is observed for XMM-Newton, though it includes a population of very soft sources lying below the X/O boundaries, potentially corresponding to starburst galaxies. For Chandra, the general trend also holds; however, there is an additional presence of very hard sources overlapping with very soft ones below the X/O boundary, particularly in the bright optical and faint X-ray regime. The majority of these outliers are sources with S/N 3 in the soft X-ray band, suggesting that at least some of them may be spurious detections. It is also plausible that these sources are intrinsically variable, suggesting that they were observed in different states of appearance. Overall, our distribution follows the findings of Merloni et al. (2014), which show that only approximately 80% of AGN selected by their SED, spectral features and HR, match their expected type I/II class.
@@ 8 Release of the catalogue
The catalogue containing the CTPs to the point-like X-ray sources in the EDFs and their properties are available through CDS/Vizier101010https://vizier.cds.unistra.fr. A detailed description of the columns and their content is provided in Appendix B. The catalogue includes unique identifiers for all relevant surveys as well as the basic X-ray and Euclid properties (columns 1 – 13). For the full list of available columns, please refer to the original catalogues introduced in Sect. 3 and e.g., Euclid Collaboration: McCracken et al. (2025); Euclid Collaboration: Polenta et al. (2025); Euclid Collaboration: Romelli et al. (2025). Columns 14 – 24 present the results of the CTP association, followed by key parameters derived from NWAY. Photometric data from LS10 are reported in columns 25 – 28. Spectroscopic information, when available, is included in columns 29 – 33. Lastly, the photometric redshift parameters from the PICZL algorithm, as well as quantities computed with such, are listed in columns 34 – 40.
@@ 9 Conclusions and outlook
In this paper we have presented the procedure adopted for the identification, classification, and study of the Euclid CTPs to the point-like sources selected from XMM-Newton, Chandra, and eROSITA X-ray, detected in Q1. We summarise the most important results in the following.
Utilising NWAY (Salvato et al. 2018), supplemented by a prior on X-ray emission, based on properties of a 4XMM DR11, and CSC 2.0 training sample, we achieved highly accurate CTP identifications relying on Euclid photometry only (see Sect. 4). Randomising the X-ray position and repeating the same procedure for the association, we quantify for each X-ray source the probability of chance associations. Users will be able to create subsamples of CTPs of desired purity and completeness by applying more restrictive or inclusive p_any (the probability of an X-ray source having a CTP) cuts (see Fig. 7).
Altogether, we identify 12 645 CTPs, with only about 10% of the sources having more than a single possible CTP, mostly in the 4XMM and eROSITA surveys, due to their larger positional uncertainties, compared to Chandra. We attribute this performance to the availability of extensive X-ray detected samples with secure CTPs, and deep, homogenised multiwavelength photometry spanning optical to MIR wavelengths, enabling robust SED construction for training. Further enhancements in expanding to broader multiwavelength coverage will improve the quality of CTP identification, particularly with Euclid DR1 in mind.
We trained a RF algorithm to classify objects as Galactic or extragalactic using solely Euclid photometry in Sect. 5. We previously trained on a sample of secure Galactic/extragalactic sources detected in LS10, using a over W1 thresholds, as presented in Salvato et al. (2022), effectively breaking the degeneracy in classification that arises when relying on Euclid photometry alone (see Euclid Collaboration: Bisigello et al. 2024). We validate these classifications by comparing to Euclid Collaboration: Matamoro Zatarain et al. (2025).
We also added spectroscopic redshifts for 16% of the CTP sample using publicly available data. For the 76% of the sample detected in LS10, we computed photometric redshifts using PICZL (Roster et al. 2024), reaching up to , although at these high redshifts the values are associated with large uncertainties. With all these redshifts, we investigated the colour-redshift relation of our sample in addition to computing their X-ray luminosities (refer to Sect. 6). We find that almost all of our CTPs pass the threshold to be considered an AGN, with most of them appearing as soft, bright, and unobscured type I sources.
@@
The RF models used for the identification of the CTPs and their classification into Galatic/extragalactic, can be applied blindly to the entire Q1 catalogue. For each source, they provide the probability of being an X-ray emitter, as well as an extragalactic source, despite the fact of not having been detected by any of the three X-ray surveys yet. Consequently, as illustrated in Figure Fig. 19, limiting the selection of Q1 sources to those with and , while excluding Euclid sources classified as CTPs in this study, yields a final sample of 135 303 objects. Notably, approximately 400 of these have been successfully cross-matched geometrically with the little red dots sample introduced by Euclid Collaboration: Bisigello et al. (2025), predominantly occupying the bright end of the magnitude distribution. A forthcoming paper will explore the multi-wavelength nature of likely extragalactic high sources that remain undetected in current X-ray surveys.
@@ Acknowledgements.
- Cuillandre et al. (2024) Cuillandre, J.-C., Bolzonella, M., Boselli, A., et al. 2024, A&A, accepted, arXiv:2405.13501
@@
- Euclid Collaboration: Aussel et al. (2025) Euclid Collaboration: Aussel, H., Tereno, I., Schirmer, M., et al. 2025, A&A, submitted
@@
- Euclid Collaboration: Bisigello et al. (2025) Euclid Collaboration: Bisigello, L., Rodighiero, G., Fotopoulou, S., et al. 2025, A&A, submitted
- Euclid Collaboration: Cropper et al. (2024) Euclid Collaboration: Cropper, M., Al Bahlawan, A., Amiaux, J., et al. 2024, A&A, accepted, arXiv:2405.13492
@@
- Euclid Collaboration: Jahnke et al. (2024) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2024, A&A, accepted, arXiv:2405.13493
@@
- Euclid Collaboration: Matamoro Zatarain et al. (2025) Euclid Collaboration: Matamoro Zatarain, T., Fotopoulou, S., Ricci, F., et al. 2025, A&A, submitted
- Euclid Collaboration: McCracken et al. (2025) Euclid Collaboration: McCracken, H., Benson, K., et al. 2025, A&A, submitted
- Euclid Collaboration: Mellier et al. (2024) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2024, A&A, accepted, arXiv:2405.13491
- Euclid Collaboration: Polenta et al. (2025) Euclid Collaboration: Polenta, G., Frailis, M., Alavi, A., et al. 2025, A&A, submitted
- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, submitted
@@
- Maciaszek et al. (2022) Maciaszek, T., Ealet, A., Gillard, W., et al. 2022, in SPIE, Vol. 12180, Space Telescopes and Instrumentation 2022: Optical, Infrared, and Millimeter Wave, ed. L. E. Coyle, S. Matsuura, & M. D. Perrin, arXiv:2210.10112
@@
- Salvato et al. (2008) Salvato, M., Hasinger, G., Ilbert, O., et al. 2008, ApJ, 690, 1250
@@ Appendix A RF parameters
Understanding which features are most relevant to our model estimates is crucial for predictive accuracy. An ML model operating as a black box may produce predictions, but the lack of transparency raises questions about the reliability of the outcomes. This not only hinders immediate interpretability but also complicates post-analyses. The ML branch investigating the role each feature plays in the model’s predictions is commonly referred to as ’feature importance’. To this end, we compute SHapley Additive exPlanations (SHAP; Lundberg & Lee 2017; Rozemberczki et al. 2022; Chen et al. 2022) values, based on a game-theoretic approach, by calculating the partial relevance for each feature. Based on a conditional expectation function, features are assigned importance values given their impact on test sample predictions. For a total achievable value , SHAP values are computed via
@@
Appendix B Column description of released CTP catalogue
@@
9.
Fx_hard: Mean hard (4XMM DR14: 2 – 4.5 keV, CSC 2.0: 2 – 7 keV and eROSITA DR1: 2–5 keV) X-ray flux from all detections of the X-ray source ()
@@
14.
Xray_EUCLID_ID: Concatenation of X-ray_ID and EUCLID_ID (unique)
15.
EUCLID_Xray_proba: Probability to be an X-ray emitter based on the RF
@@
18.
dist_post: Distance posterior probability comparing this association vs. no association (see Appendix in Salvato et al. 2018)
19.
@@
28.
Gal_proba: Probability of being a Galactic rather than extragalactic source based on RF
@@
Cat_spec: Parent catalogue of the spec- match
@@
phz_1sl: Photo- lower 1 sigma error
@@
phz_1su: Photo- upper 1 sigma error
@@
phz_3sl: Photo- lower 3 sigma error
@@
phz_3su: Photo- upper 3 sigma error
@@
z_final: Spec- when available; otherwise, photo-
40.
Lx: X-ray luminosity, assuming FlatLambdaCDM from astropy.cosmology

### v2

@@ 7.3 AGN type
In Fig. 18, we plot the respective -band magnitudes as a function of the soft X-ray fluxes for all CTPs classified as extragalactic. We add the —X/O— = 1 lines (Maccacaro et al. 1988), which confine the typical AGN locus (see also e.g., Salvato et al. 2011; Civano et al. 2012), defined as
where is the X-ray flux in a given energy range, and correspond to the magnitude and a constant which both depend on the specific filter used in the optical observations. For we find . To further assess the AGN type of our CTPs, we colour-code the plots in the bottom row of Fig. 18 by the X-ray luminosity and the top row by hardness ratio (Hasinger et al. 2005; Cappelluti et al. 2007; Salvato et al. 2009), defined as
where soft corresponds to the 0.5–2.0 keV detection band, and hard refers to 2.0–12.0 keV PN thin for 4XMM DR14, 2.0–7.0 keV ACIS-I for Chandra, and 2.0–5.0 keV for eROSITA DR1 respectively, to account for each instrument’s flux sensitivity. This ensures as much consistency as possible across the different surveys, to make the different instruments somewhat comparable. A negative HR value indicates soft, unobscured sources (type I), and a positive HR value suggests harder, likely obscured sources (type II). We examine the HR distributions for all three telescopes to assess differences in their observed AGN populations. However, a non-negligible fraction of sources lack hard-band detections: approximately 47.6% for eROSITA, 14.5% for XMM-Newton, and 7.3% for . These non-detections, especially in the case of eROSITA, naturally shift the medians of the true HR distributions and introduce selection biases. As shown in Fig. 24, we find that the eROSITA HR distribution median is , while that of and (XMM-Newton) show a similar tendency towards harder values with medians of . These HR medians can then be used to compute ‘effective’ photon indices , which we find to be for eROSITA, for 131313Adopting CHANDRA-Cycle 11 as a representative benchmark, as it reflects an intermediate instrumental state within the 20-year span of observations included in CSC 2.0, accounting for the effects of detector degradation over time. and for XMM-Newton, assuming negligible column density. As expected, the greater sensitivity of XMM-Newton and at higher energies allows them to detect harder and possibly mildly obscured AGNs that are missed by eROSITA. Conversely, eROSITA tends to detect brighter AGN which exhibit softer X-ray spectra. As the markers in Fig. 18 are plotted in order of increasing HR, many softer sources in the background are being covered up and predominantly fall within the X/O boundaries, classifying them as AGN, with differently weighted populations visible for every X-ray survey. As observed for XMM-Newton and , this includes a population of sources with lying below the X/O boundaries, potentially corresponding to starburst galaxies, low-luminosity AGN, or obscured AGN. However, the majority of these sources show S/Nx 3 in the soft X-ray band, suggesting that a few of these may be spurious detections. It is also plausible that these sources are intrinsically variable, suggesting that they were observed in different states of appearance. Overall, our distribution follows the findings of Merloni et al. (2014), which show that only approximately 80% of AGN selected by their SED, spectral features and HR, match their expected type I/II class.
@@ 8 Release of the catalogue
The catalogue containing the CTP properties of the EDF point-like X-ray sources is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/. A detailed description of the columns and their content is provided in Appendix D. The catalogue includes unique identifiers for all relevant surveys as well as the basic X-ray and Euclid properties (columns 1–14). For the full list of available columns, please refer to the original catalogues introduced in Sect. 3 and, for example, Euclid Collaboration: McCracken et al. (2025); Euclid Collaboration: Polenta et al. (2025); Euclid Collaboration: Romelli et al. (2025). Columns 15–24 present the results of the CTP association, followed by key parameters derived from NWAY. Photometric data from LS10 are reported in columns 25–27. Spectroscopic information, when available, is included in columns 28–32. Lastly, the photometric redshift parameters from the PICZL algorithm, as well as quantities computed with such, are listed in columns 33–39.
@@ 9 Conclusions and outlook
In this paper, we have presented the procedure adopted for the identification, classification, and study of the Euclid CTPs of the point-like sources selected from XMM-Newton, Chandra, and eROSITA X-ray, detected in Q1. We summarise the most important results in the following.
Utilising NWAY (Salvato et al. 2018) supplemented by a prior on X-ray emission based on properties of a 4XMM DR11 and CSC 2.0 training sample, we achieved highly accurate CTP identifications while relying on Euclid photometry only (see Sect. 4). By randomising the X-ray position and repeating the same procedure for the association, we quantified for each X-ray source the probability of chance associations. Users will be able to create subsamples of CTPs of desired purity and completeness by applying more restrictive or inclusive p_any cuts ( probability of an X-ray source having a CTP; see Fig. 7).
Altogether, we have identified 12 645 CTPs, with only about 10% of the sources having more than a single possible CTP. These are mostly in the XMM-Newton and eROSITA surveys, due to their larger positional uncertainties compared to Chandra. We attribute this performance to the availability of extensive X-ray detected samples with secure CTPs and deep, homogenised multiwavelength photometry spanning optical to MIR wavelengths enabling robust SED construction for training. Further expansion to broader multiwavelength coverage will improve the quality of CTP identification, particularly for Euclid DR1.
We trained an RF algorithm to classify objects as Galactic or extragalactic using solely Euclid photometry in Sect. 5. We previously trained another RF on a sample of secure Galactic and extragalactic sources detected in LS10, using over W1 thresholds, as presented in Salvato et al. (2022), effectively breaking the degeneracy in classification that arises when relying on Euclid photometry alone (see Euclid Collaboration: Bisigello et al. 2024). We validated these classifications by comparing to Euclid Collaboration: Matamoro Zatarain et al. (2025).
We also added spectroscopic redshifts for 16% of the CTP sample using publicly available data. For the 82% of the sample detected in LS10, we computed photo- using PICZL (Roster et al. 2024), reaching up to . However at these high redshifts, the values are associated with large uncertainties. With all of these redshifts, we investigated the colour-redshift relation of our sample in addition to computing their X-ray luminosities (refer to Sect. 6). We find that all of our CTPs pass our soft X-ray luminosity threshold to be considered an AGN, with more of them appearing as bright type I sources rather than type II.
@@
The RF models used for the identification of the CTPs and their classification as Galactic or extragalactic, can be applied blindly to the entire Q1 catalogue. For each source, they provide the probability of being an X-ray emitter as well as an extragalactic source, despite not having been detected by any of the three X-ray surveys yet. Consequently, as illustrated in Fig. 19, limiting the selection of Q1 sources to those with and while excluding Euclid sources classified as CTPs in this study yields a final sample of 135 303 objects. Notably, 73 of these objects have been successfully cross-matched geometrically with the little red dots (LRDs) sample introduced by Euclid Collaboration: Bisigello et al. (2025), where the matches predominantly occupy the bright end of the LRD magnitude distribution. A forthcoming paper will explore the multi-wavelength nature of the likely extragalactic high sources that remain undetected in current X-ray surveys.
@@ Acknowledgements.
- Cuillandre et al. (2025) Cuillandre, J.-C., Bolzonella, M., Boselli, A., et al. 2025, A&A, 697, A11
@@
- Euclid Collaboration: Aussel et al. (2025) Euclid Collaboration: Aussel, H., Tereno, I., Schirmer, M., et al. 2025, A&A, submitted (Euclid Q1 SI), arXiv:2503.15302
@@
- Euclid Collaboration: Bisigello et al. (2025) Euclid Collaboration: Bisigello, L., Rodighiero, G., Fotopoulou, S., et al. 2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554537, arXiv:2503.15323
- Euclid Collaboration: Cropper et al. (2025) Euclid Collaboration: Cropper, M., Al-Bahlawan, A., Amiaux, J., et al. 2025, A&A, 697, A2
@@
- Euclid Collaboration: Jahnke et al. (2025) Euclid Collaboration: Jahnke, K., Gillard, W., Schirmer, M., et al. 2025, A&A, 697, A3
@@
- Euclid Collaboration: Matamoro Zatarain et al. (2025) Euclid Collaboration: Matamoro Zatarain, T., Fotopoulou, S., Ricci, F., et al. 2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554619, arXiv:2503.15320
- Euclid Collaboration: McCracken et al. (2025) Euclid Collaboration: McCracken, H. J., Benson, K., Dolding, C., et al. 2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554594, arXiv:2503.15303
- Euclid Collaboration: Mellier et al. (2025) Euclid Collaboration: Mellier, Y., Abdurro’uf, Acevedo Barroso, J., et al. 2025, A&A, 697, A1
- Euclid Collaboration: Polenta et al. (2025) Euclid Collaboration: Polenta, G., Frailis, M., Alavi, A., et al. 2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554657, arXiv:2503.15304
- Euclid Collaboration: Romelli et al. (2025) Euclid Collaboration: Romelli, E., Kümmel, M., Dole, H., et al. 2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554586, arXiv:2503.15305
@@
- Euclid Collaboration: Tucci et al. (2025) Euclid Collaboration: Tucci, M., Paltani, S., Hartley, W. G., et al. 2025, A&A, in press (Euclid Q1 SI), https://doi.org/10.1051/0004-6361/202554588, arXiv:2503.15306
@@
- Giavalisco et al. (2004) Giavalisco, M., Ferguson, H. C., Koekemoer, A. M., et al. 2004, ApJ, 600, L93
- Graham et al. (2020) Graham, M. L., Connolly, A. J., Wang, W., et al. 2020, AJ, 159, 258
@@
- Maciaszek et al. (2022) Maciaszek, T., Ealet, A., Gillard, W., et al. 2022, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 12180, Space Telescopes and Instrumentation 2022: Optical, Infrared, and Millimeter Wave, ed. L. E. Coyle, S. Matsuura, & M. D. Perrin, arXiv:2210.10112
@@
- Martin et al. (2005) Martin, D. C., Fanson, J., Schiminovich, D., et al. 2005, ApJ, 619, L1
- McMahon et al. (2013) McMahon, R. G., Banerji, M., Gonzalez, E., et al. 2013, The Messenger, 154, 35
@@
- Salvato et al. (2009) Salvato, M., Hasinger, G., Ilbert, O., et al. 2009, ApJ, 690, 1250
@@
- Zhou et al. (2023) Zhou, R., Ferraro, S., White, M., et al. 2023, J. Cosmology Astropart. Phys., 2023, 097
- Zhou et al. (2021) Zhou, R., Newman, J. A., Mao, Y.-Y., et al. 2021, MNRAS, 501, 3309
@@ Appendix A RF parameters
Understanding which features are most relevant to our model estimates is crucial for predictive accuracy. An ML model operating as a black box may produce predictions, but the lack of transparency raises questions about the reliability of the outcomes. This not only hinders immediate interpretability but also complicates post-analyses. The ML branch investigating the role each feature plays in the model’s predictions is commonly referred to as ‘feature importance’. To this end, we compute SHapley Additive exPlanations (SHAP; Lundberg & Lee 2017; Rozemberczki et al. 2022; Chen et al. 2022) values, based on a game-theoretic approach, by calculating the partial relevance for each feature. Based on a conditional expectation function, features are assigned importance values given their impact on test sample predictions. For a total achievable value , SHAP values are computed via
@@
Appendix B Photometric redshift comparison
Appendix C HR to effective photon index
Appendix D Column description of released CTP catalogue
@@
HR: Count rate derived hardness ratio considering soft and hard (4XMM DR14: , CSC 2.0: , and eROSITA: ) bands.
11.
EDF: Euclid deep field
12.
@@
EUCLID_P_Xray: Probability to be an X-ray emitter based on the RF
17.
EUCLID_P_Gal: Probability of being a Galactic rather than extragalactic source based on the RF
18.
@@
20.
dist_post: Distance posterior probability comparing this association versus no association (see Appendix in Salvato et al. 2018)
21.
@@
Cat_spec: Parent catalogue of the spec-
@@
phz_1sl: PICZL lower 1 sigma error
@@
phz_1su: PICZL upper 1 sigma error
@@
phz_Luo17: Photometric redshift (see Luo et al. 2017)
@@
z_final: Spec- when available; otherwise Luo et al. (2017) and PICZL photo-
@@
Lx: Rest-frame X-ray luminosity, assuming FlatLambdaCDM from astropy.cosmology